Find the coordinates of the midpoint of a segment with each pair of endpoints.
K(1/4, 1), L(1/8, 0)

Answer by checkley71(8403)

MIDPOINTS ARE FOR X (1/4+1/8)/2 OR ((2+1)/8)/2=(3/8)/2=3/16 THUS X MIDPOINT IS 3/16 & THE Y MID POINT IS 0+1/2 OR 1/2 THUS THE MIDPOINT IS (3/16,1/2)